OpenStudy (anonymous):

Marco is putting some sweaters into storage for the winter. He has 10 sweaters, and he can fit 6 sweaters into a box. How many different groups of 6 sweaters can Marco pack into a box? (1 point)

OpenStudy (kropot72):

\[10C6=\frac{10!}{(10-6)!6!}=\frac{10\times 9\times 8\times 7}{4\times 3\times 2\times 1}\]

OpenStudy (anonymous):

I got its 210 for the answer @kropot72

OpenStudy (kropot72):

Good work! There are 210 different groups.

OpenStudy (anonymous):

thanks ;)

OpenStudy (kropot72):

You're welcome :)